# Akchen and Mišić

### v3

@@ 1 Introduction
In the present paper, we present a methodology for assortment optimization under the decision forest model, based on mixed-integer optimization. Our approach allows the firm to obtain assortments that are either provably optimal or within a desired optimality gap for a given decision forest model. At the same time, the approach easily allows the firm to incorporate business rules as linear constraints in the MIO formulation. Most importantly, given the universality property of the decision forest model, our optimization approach allows a firm to optimally tailor its assortment to any kind of predictable irrationality in the firm’s customer population.
@@
    We formally define the assortment optimization problem under the decision forest model and analyze its computational hardness. Specifically, assuming the tree depth is bounded by a constant $k$, we prove that it is NP-hard to approximate the problem within a factor that grows exponentially with $k$. Given this computational challenge, we then adopt an integer programming approach and present two formulations: SplitMIO and ProductMIO. We analyze the constraint matrices of the two formulations and show that in the special case of a single purchase decision tree, SplitMIO is integral when each product appears at most once in the splits of the tree, and ProductMIO is always integral regardless of the tree structure.
@@
    We propose a Benders decomposition approach for solving the two formulations at a large scale. We show that Benders cuts for the linear optimization relaxations of SplitMIO can be obtained via a greedy algorithm that solves both the primal and dual of the subproblem. We also provide a simple example to show that the same type of greedy algorithm fails to solve the primal subproblem of ProductMIO. We further show how to obtain Benders cuts for the integer solutions of both SplitMIO and ProductMIO in closed form.
@@
    We present numerical experiments using both synthetic and real-world data. We first use the synthetic data to examine the formulation strength of SplitMIO and ProductMIO models, and to demonstrate the scalability of the Benders decomposition approach for the SplitMIO formulation in problem instances involving up to 3000 products, 500 trees, and 512 leaves per tree. We then use a real-world dataset to demonstrate how the decision forest model can factor a behavioral anomaly (choice overload) into its assortment decision and create higher revenue.
@@ 2 Literature review
Assortment optimization has been extensively studied in the operations management community; we refer readers to Gallego and Topaloglu (2019) for a recent review of the literature. The literature on assortment optimization has focused on developing approaches for finding the optimal assortment under many different rational choice models, such as the multinomial logit model (MNL) model (Talluri and Van Ryzin 2004, Sumida et al. 2020), the latent class MNL model (Bront et al. 2009, Méndez-Díaz et al. 2014), the Markov chain choice model (Feldman and Topaloglu 2017, Désir et al. 2020) and the ranking-based model (Aouad et al. 2020, 2018, Feldman et al. 2019).
@@
In addition to the assortment optimization and product line design literatures, our formulations also have connections with others that have been proposed in the broader optimization literature. The formulation SplitMIO that we will present later can be viewed as a special case of the mixed-integer optimization formulation of Mišić (2020) for optimizing the predicted value of a tree ensemble model, such as a random forest or a boosted tree model. The other formulation, ProductMIO, also has a connection to the integer optimization literature on formulating disjunctive constraints through independent branching schemes (Vielma et al. 2010, Vielma and Nemhauser 2011, Huchette and Vielma 2019); we also discuss this connection in more detail in Section 3.4.
@@ 3 Optimization model
In this section, we define the decision forest assortment optimization problem (Section 3.1), analyze its computational complexity (Section 3.2), and subsequently present the two mixed-integer formulations, SplitMIO (Section 3.3) and ProductMIO (Section 3.4). Lastly, we show that when the number of trees is assumed to be constant, the decision forest assortment optimization problem can be solved in polynomial time (Section 3.5).
@@ 3.1 Problem definition
We first briefly review the decision forest model of Chen and Mišić (2022), and then formally state the assortment optimization problem. We assume that there are $n$ products, indexed from 1 to $n$, and let $\mathcal{N}=\{1,\dots,n\}$ denote the set of all products. An assortment $S$ corresponds to a subset of $\mathcal{N}$. When offered $S$, a customer may choose to purchase one of the products in $S$, or to not purchase anything from $S$ at all; we use the index 00 to denote the latter possibility, which we will also refer to as the no-purchase option or the outside option.
@@
Figure 1: (left) Example of a purchase decision tree for $n=5$ products. Leaf nodes are enclosed in squares, while split nodes are not enclosed. The number on each node corresponds either to $v(t,s)$ for splits, or $c(t,\ell)$ for leaves. The path highlighted in red indicates how a customer following this tree to maps the assortment $S=\{1,3,4,5\}$ to a leaf. For this assortment, the customer’s decision is to purchase product 5; (right) Example of a purchase decision tree that represents a preference ranking $\{3\succ 1\succ 2\succ 0\}$.
@@
$\mathbf{P}^{(F,\boldsymbol{\lambda})}(j\mid S)=\sum_{t\in F}\lambda_{t}\cdot\mathbb{I}\{{\color[rgb]{0,0,0}\eta(t,S)}=j\}.$ (1)
@@ 3.2 Inapproximability Results
If P $\neq$ NP, for every constant $k\geq 7$ and $\epsilon>0$, there is no polynomial-time $\left(2^{k-2\lceil\sqrt{k}\rceil}/(2ek)-\epsilon\right)$-approximate algorithm for AODF(k+1), even when each tree has equal probability weight in the model, i.e., $\lambda_{t}=1/|F|$.
[…]
@@
Before we proceed to solve the challenging assortment optimization problem (2), we pause to make two remarks on Theorem 3.2. First, recall that any ranking of length $k$ can be represented by a decision tree that is constrained to branch to the right and of depth $k$. Aouad et al. (2018) shows that the assortment optimization under the ranking-based model is NP-hard to be approximated within $O(k^{1-\epsilon})$ factor provided that each ranking in the model has a length of at most $k$ (see Section 3.2 of Aouad et al. (2018)). Theorem 3.2 shows that the assortment problem becomes much harder when optimizing over decision trees instead of rankings. Specifically, the inapproximability factor grows from a linear rate $O(k)$ to an exponential rate $O(2^{k})$ as a function of $k$.
Second, together with Chen and Mišić (2022), Theorem 3.2 highlights a fundamental tradeoff between the representation power and tractability in the decision forest model. In particular, Chen and Mišić (2022) shows that as the decision trees grow deeper, the model becomes more expressive in capturing choice behavior; at its maximum depth, $n$, the decision forest model can represent any discrete choice model (see Theorems 1 and 2 in Chen and Mišić (2022)). At the same time, our Theorem 3.2 establishes that deeper trees also lead to increased computational complexity for assortment optimization, with inapproximability growing exponentially in tree depth. This tradeoff underscores the importance of managing tree depth to ensure that the model accurately reflects consumer choice while maintaining tractability for downstream operational tasks.
@@ 3.3 Formulation 1: SplitMIO
[…]
@@
In order of appearance, the constraints in this formulation have the following meaning. Constraint (3b) requires that for each customer type $t$, the assortment encoded by $\mathbf{x}$ is mapped to exactly one leaf. Constraint (3c) imposes that for a split $s$ in tree $t$, if product $v(t,s)$ is not in the assortment, then the assortment cannot be mapped to any of the leaves that are to the left of split $s$ in tree $t$. Constraint (3d) is the symmetric case of constraint (3c) for the right-hand subtree of split $s$ of tree $t$. The last two constraints require that $\mathbf{x}$ is binary and $\mathbf{y}$ is nonnegative. Note that it is not necessary to require $\mathbf{y}$ to be binary, as the constraints ensure that each $y_{t,\ell}$ automatically takes the correct value whenever $\mathbf{x}$ is binary. Finally, the objective function corresponds to the expected per-customer revenue of the assortment.
@@
Before continuing to our second formulation, we establish an important property of problem (3) when $|F|=1$. When $|F|=1$, we can show that $\mathcal{F}_{\textsc{SplitMIO}}$ is integral in a particular special case. (Note that in the statement of the proposition below, we drop the index $t$ to simplify notation.)
Proposition 3.4
Let $(F,\boldsymbol{\lambda})$ be a decision forest model consisting of a single tree, i.e., $|F|=1$. Then $\mathcal{F}_{\textsc{SplitMIO}}$ is integral, i.e., every extreme point $(\mathbf{x},\mathbf{y})$ of the polyhedron $\mathcal{F}_{\textsc{SplitMIO}}$ satisfies $\mathbf{x}\in\{0,1\}^{N}$, if and only if $v(s)=i$ for at most one $s\in\mathbf{splits}$ for every product $i\in\mathcal{N}$.
The proof of this result (see Appendix 8.3) follows by showing that the constraint matrix defining $\mathcal{F}_{\textsc{SplitMIO}}$ is totally unimodular. Proposition 3.4 is significant because it implies that for $|F|=1$, the distinction between trees where each product appears at most once in any split and trees where a product may appear two or more times as a split is sharp. This insight provides the motivation for our second formulation, ProductMIO, which we present next.
@@ 3.4 Formulation 2: ProductMIO
[…]
@@
ProductMIO differs from SplitMIO in several ways. First, note that while both formulations have the same number of variables, formulation ProductMIO has a smaller number of constraints. In particular, SplitMIO has one left and one right split constraints for each split in each tree, whereas ProductMIO has one left and one right split constraint for each product. When the trees involve a large number of splits, this can lead to a sizable reduction in the number of constraints. Note also that when a product does not appear in any splits of a tree, we can also safely omit constraints (7c) and (7d) for that product.
The second difference with formulation SplitMIO, as we have already mentioned, is in formulation strength. Let $\mathcal{F}_{\textsc{ProductMIO}}$ be the feasible region of the linear optimization (LO) relaxation of ProductMIO. The following proposition formalizes the fact that formulation ProductMIO is at least as strong as formulation SplitMIO.
@@
The last major difference is in how ProductMIO behaves when $|F|=1$. We saw that a sufficient condition for $\mathcal{F}_{\textsc{SplitMIO}}$ to be integral when $|F|=1$ is that each product appears in at most one split in the tree. In contrast, formulation ProductMIO is always integral when $|F|=1$.
Proposition 3.6
For any decision forest model $(F,\boldsymbol{\lambda})$ with $|F|=1$, $\mathcal{F}_{\textsc{ProductMIO}}$ is integral.
In contrast to Proposition 3.4, we take a different approach to proving the integrality property, as the constraint matrix of the ProductMIO formulation is generally not totally unimodular (see the example in Section 8.4). The proof of Proposition 3.6, given in Appendix 8.5, leverages a connection between ProductMIO and another type of formulation in the literature. In particular, a stream of papers in the mixed-integer optimization community (Vielma et al. 2010, Vielma and Nemhauser 2011, Huchette and Vielma 2019) has considered a general approach for deriving small and strong formulations of disjunctive constraints using independent branching schemes; we briefly review the most general such approach from Huchette and Vielma (2019) and showcase its connection to ProductMIO. In this approach, one has a finite ground set $J$, and is interested in optimizing over a particular subset of the $(|J|-1)-$dimensional unit simplex over $J$, $\Delta^{J}=\{\boldsymbol{\lambda}\in\mathbb{R}^{J}\mid\sum_{j\in J}\lambda_{j}=1;\boldsymbol{\lambda}\geq\mathbf{0}\}$. The specific subset that we are interested in is called a combinatorial disjunctive constraint (CDC), and is given by $\text{CDC}(\mathcal{S})=\bigcup_{S\in\mathcal{S}}Q(S)$, where $\mathcal{S}$ is a finite collection of subsets of $J$ and $Q(S)=\{\boldsymbol{\lambda}\in\Delta\mid\lambda_{j}\leq 0\ \text{for}\ j\in J\setminus S\}$ for any $S\subseteq J$. This approach is very general: for example, by associating each $j$ with a point $\mathbf{x}^{j}$ in $\mathbb{R}^{n}$, one can use $\text{CDC}(\mathcal{S})$ to model an optimization problem over a union of polyhedra, where each polyhedron is the convex hull of a collection of vertices in $S\in\mathcal{S}$.
A $k$-way independent branching scheme of depth $t$ is a representation of $\text{CDC}(\mathcal{S})$ as a sequence of $t$ choices between $k$ alternatives. Specifically, we have $\text{CDC}(\mathcal{S})=\bigcap_{m=1}^{t}\bigcup_{i=1}^{k}Q({\color[rgb]{0,0,0}L^{i}_{m}})$, where ${\color[rgb]{0,0,0}L^{i}_{m}}\subseteq J$. In the special case that $k=2$, we can write $\text{CDC}(\mathcal{S})=\cap_{m=1}^{t}(L_{{m}}\cup R_{{m}})$ where $L_{{m}},R_{{m}}\subseteq J$. This representation is known as a pairwise independent branching scheme and the constraints of the corresponding MIO can be written simply as
@@
This particular special case is important because it is always integral (see Theorem 1 of Vielma et al. 2010). Moreover, we can see that ProductMIO bears a strong resemblance to formulation (8). Constraints (8a) and (8a) correspond to constraints (7c) and (7d), respectively. In terms of variables, the $\lambda_{j}$ and $z_{m}$ variables in formulation (8) correspond to the $y_{t,\ell}$ and $x_{i}$ variables in ProductMIO, respectively.
@@
3.5 Polynomial time algorithm when $|F|=O(1)$
We conclude Section 3 by discussing a special setting in which the assortment problem (2) is efficiently solvable. When $|F|=1$, the assortment optimization problem is trivial to solve, because one can simply find which leaf corresponds to the highest value of $r_{t,\ell}$, and then follow the path from the leaf to the root backwards to identify which products must be present and absent in the assortment to ensure that it is mapped to this leaf. If the number of leaves is polynomial in $n$, then one can solve this problem in time polynomial in $n$.
In the case that $|F|=O(1)$, i.e., the number of trees is assumed to be constant, it turns out that the assortment problem can also be solved efficiently, which we formalize in the proposition below.
Proposition 3.7
Assume $|F|=O(1)$ and $|\mathbf{leaves}(t)|$ is polynomial in $n$ for each tree $t\in F$. Then the assortment problem (2) can be solved optimally in a polynomial time.
Combined with Theorem 3.2, Proposition 3.7 shows that it is the number of trees $|F|$, rather than the structure of each tree (e.g., the number of leaves) that drives the complexity of the assortment problem (2) from being tractable to inapproximable (unless P = NP). Recall that in the reduction that proves Theorem 3.2, the number of leaf nodes in each tree is roughly $k^{2}/2$, which is $O(1)$ under the theorem’s assumption of constant $k$. Consequently, compared to Theorem 3.2, Proposition 3.7 suggests that even if the trees become significantly more complex – where the number of leaves in each tree increases from $O(1)$ to a polynomial in $n$ – the assortment optimization problem remains solvable in polynomial time, provided that the number of trees $|F|$ is $O(1)$.

### v4

@@ 1 Introduction
In the present paper, we present a methodology for assortment optimization under the decision forest model, based on mixed-integer optimization. Our approach allows the firm to obtain assortments that are either provably optimal or within a desired optimality gap for a given decision forest model. Moreover, in general, business rules (e.g., capacity constraints) can be readily incorporated by adding corresponding linear constraints to the MIO formulation. Most importantly, given the universality property of the decision forest model, our optimization approach allows a firm to optimally tailor its assortment to any kind of predictable irrationality in the firm’s customer population.
@@
    We formally define the assortment optimization problem under the decision forest model and analyze its computational hardness. Specifically, when the tree depth is bounded by a constant $k$, we prove that it is NP-hard to approximate the problem within a factor that grows exponentially in $k$. This result complements the findings of Chen and Mišić (2022), who show that tree depth governs the representational power of the model: with sufficiently deep trees, the decision forest can represent any discrete choice model. However, they also note that deeper trees increase the risk of overfitting and recommend selecting tree depth via cross-validation. From an operations perspective, our hardness result further underscores the importance of shallow trees: while deeper trees enhance representation power, they also render the downstream assortment optimization problem inapproximable, highlighting the need to balance model expressiveness with tractability. Given the computational challenge of the assortment optimization problem under the decision forest model, we then adopt an integer programming approach and present two formulations: SplitMIO and ProductMIO. We compare the strength of the two formulations and provide the integrality analysis in the single-tree case.
@@
    We propose a Benders decomposition approach for solving the two formulations at scale. We apply the method to both the linearly relaxed and the integer programs so that the cuts generated from the relaxation provide strong upper bounds and warm-start the integer procedure. For SplitMIO, we show that Benders cuts for the linear relaxation can be obtained via a greedy algorithm that simultaneously solves the primal and dual subproblems. We also give a simple example demonstrating that this greedy construction fails for the primal subproblem of ProductMIO. For both SplitMIO and ProductMIO, we further derive closed-form Benders cuts for the integer subproblems. Overall, our decomposition generalizes the scheme developed for the ranking-based model (Bertsimas and Mišić 2019) by introducing a new dual-variable update procedure that exploits the tree topology and is tailored to the subproblem structures of our formulations.
@@
    We present three numerical experiments using both synthetic and real-world data. The first, based on synthetic instances, evaluates the formulation strength of SplitMIO and ProductMIO and demonstrates the scalability of the Benders decomposition approach for SplitMIO on problems with up to 3,000 products, 500 trees, and 512 leaves per tree. The second uses the IRI Dataset (Bronnenberg et al. 2008), a large-scale grocery transaction dataset, to show how the decision forest model can capture a behavioral anomaly (choice overload) and recommend assortments that differ from those implied by rational choice models. The third is a semi-synthetic, end-to-end (data-to-decision) experiment: we generate data from a ranking-based model derived from the Sushi Dataset (Kamishima 2003), augmented with a choice overload effect (Iyengar and Lepper 2000). We show that the decision forest model delivers robust and consistent performance even as the overload effect intensifies, whereas the multinomial logit (MNL) and ranking-based models deteriorate noticeably, even under mild choice overload. Taken together, the second and third experiments highlight that the decision forest model can effectively incorporate behavioral anomalies observed in consumer choice into the assortment decision process.
@@ 2 Literature review
Assortment optimization has been extensively studied in the operations management community; we refer readers to Gallego and Topaloglu (2019) for a recent review of the literature. The literature on assortment optimization has focused on developing approaches for finding the optimal assortment under many different rational choice models, such as the MNL model (Talluri and Van Ryzin 2004, Sumida et al. 2020), the latent class MNL model (Bront et al. 2009, Méndez-Díaz et al. 2014), the Markov chain choice model (Feldman and Topaloglu 2017, Désir et al. 2020) and the ranking-based model (Aouad et al. 2020, 2018, Feldman et al. 2019).
@@
In addition to the assortment optimization and product line design literatures, our formulations also relate to others that have been proposed in the broader optimization literature. The formulation SplitMIO that we will present later can be viewed as a special case of the mixed-integer optimization formulation of Mišić (2020) for optimizing the predicted value of a tree ensemble model, such as a random forest or a boosted tree model. The other formulation, ProductMIO, also has a connection to the integer optimization literature on formulating disjunctive constraints through independent branching schemes (Vielma et al. 2010, Vielma and Nemhauser 2011, Huchette and Vielma 2019); we also discuss this connection in more detail in Section 3.4.
@@ 3 Optimization model
In this section, we define the decision forest assortment optimization problem (Section 3.1), analyze its computational complexity (Section 3.2), and present two mixed-integer formulations: SplitMIO (Section 3.3) and ProductMIO (Section 3.4). We then compare these formulations and highlight their respective advantages across different regimes (Section 3.5).
@@ 3.1 Problem definition
We first briefly review the decision forest model of Chen and Mišić (2022), and then formally state the assortment optimization problem. We assume that there are $n$ products, indexed from 1 to $n$, and let $\mathcal{N}=\{1,\dots,n\}$ denote the set of all products. An assortment $S$ corresponds to a subset of $\mathcal{N}$. When offered $S$, a customer may choose to purchase one of the products in $S$, or to not purchase anything from $S$ at all; we use the index 0 to denote the latter possibility, which we will also refer to as the no-purchase option or the outside option.
@@
Figure 1: (left) Example of a purchase decision tree for $n=5$ products. Leaf nodes are enclosed in squares, while split nodes are not enclosed. The number on each node corresponds either to $v(t,s)$ for splits, or $c(t,\ell)$ for leaves. The path highlighted in red shows how a customer represented by this tree makes a purchase decision when facing the assortment $S=\{1,3,4,5\}$. For this assortment, the customer’s decision is to purchase product 5; (right) Example of a purchase decision tree that represents a preference ranking $\{3\succ 1\succ 2\succ 0\}$.
@@
$\mathbf{P}^{(F,\boldsymbol{\lambda})}(j\mid S)=\sum_{t\in F}\lambda_{t}\cdot\mathbb{I}\{\eta(t,S)=j\}.$ (1)
@@ 3.2 Inapproximability Results
If P $\neq$ NP, for every constant $k\geq 7$ and $\epsilon>0$, there is no polynomial-time $\left(2^{k-2\lceil\sqrt{k}\rceil}/(2ek)-\epsilon\right)$-approximation algorithm for AODF(k+1), even when each tree has equal probability weight in the model, i.e., $\lambda_{t}=1/|F|$.
We prove Theorem 3.2 by constructing an approximation-preserving reduction between $\text{AODF}(k+1)$ and the Max 1-in-EkkSAT problem (Guruswami and Trevisan 2005), for which the same inapproximability result holds. The Max 1-in-EkkSAT problem is a variant of the classic SAT problem (Garey and Johnson 1979), where each clause is satisfied if and only if exactly one literal in the clause is true, and each clause consists of exactly $k$ literals. In our reduction, we construct a decision tree with depth $k+1$ and $O(k^{2})$ leaf nodes for each clause in the Max 1-in-EkkSAT problem.
@@
Before we proceed to solve the challenging assortment optimization problem (2), we pause to make two remarks on Theorem 3.2. First, recall that any ranking of length $k$ can be represented by a decision tree that is constrained to branch to the right and of depth $k$. Aouad et al. (2018) shows that the assortment optimization under the ranking-based model is NP-hard to approximate within $O(k^{1-\epsilon})$ factor provided that each ranking in the model has a length of at most $k$ (see Section 3.2 of Aouad et al. (2018)). Theorem 3.2 shows that the assortment problem becomes much harder when optimizing over decision trees instead of rankings. Specifically, the inapproximability factor grows from a linear rate $O(k)$ to an exponential rate $O(2^{k})$ as a function of $k$.
Second, together with Chen and Mišić (2022), Theorem 3.2 highlights a fundamental tradeoff between the representation power and tractability in the decision forest model. In particular, Chen and Mišić (2022) shows that as the decision trees grow deeper, the model becomes more expressive in capturing choice behavior; at its maximum depth, $n$, the decision forest model can represent any discrete choice model (see Theorems 1 and 2 in Chen and Mišić (2022)). At the same time, our Theorem 3.2 establishes that deeper trees also lead to increased computational complexity for assortment optimization, with inapproximability growing exponentially in tree depth. This tradeoff highlights the need to manage tree depth so that the decision forest model remains expressive enough to characterize consumer choice yet tractable for downstream operational tasks.
@@ 3.3 Formulation 1: SplitMIO
$\textsc{SplitMIO}:\quad$ $\underset{\mathbf{x},\mathbf{y}}{\text{maximize}}$ $\sum_{t\in F}\lambda_{t}\cdot\left[\sum_{\ell\in\mathbf{leaves}(t)}r_{t,\ell}y_{t,\ell}\right]$ (3a)
@@
In order of appearance, the constraints in this formulation have the following meaning. Constraint (3b) requires that for each customer type $t$, the assortment encoded by $\mathbf{x}$ is mapped to exactly one leaf. Constraint (3c) imposes that for a split $s$ in tree $t$, if product $v(t,s)$ is not in the assortment, then the assortment cannot be mapped to any of the leaves that are to the left of split $s$ in tree $t$. Constraint (3d) is the symmetric case of constraint (3c) for the right-hand subtree of split $s$ of tree $t$. The last two constraints require that $\mathbf{x}$ is binary and $\mathbf{y}$ is nonnegative. Note that it is not necessary to require $\mathbf{y}$ to be binary. Whenever $\mathbf{x}$ is binary, there exists an integer-valued solution for $\mathbf{y}$. Finally, the objective function corresponds to the expected per-customer revenue of the assortment.
@@ 3.4 Formulation 2: ProductMIO
$\textsc{ProductMIO}:\quad$ $\underset{\mathbf{x},\mathbf{y}}{\text{maximize}}$ $\sum_{t\in F}\lambda_{t}\cdot\left[\sum_{\ell\in\mathbf{leaves}(t)}r_{t,\ell}y_{t,\ell}\right]$ (7a)
@@
Note that while both formulations have the same number of variables, formulation ProductMIO has a smaller or equal number of constraints compared with SplitMIO. In particular, SplitMIO has one left and one right split constraint for each split in each tree, whereas ProductMIO has one left and one right split constraint for each product. When the trees involve a large number of splits, this can lead to a sizable reduction in the number of constraints. Additionally, when a product does not appear in any splits of a tree, we can also safely omit constraints (7c) and (7d) for that product. As we noted earlier, formulation ProductMIO is at least as strong as formulation SplitMIO. Let $\mathcal{F}_{\textsc{ProductMIO}}$ be the feasible region of the linear optimization (LO) relaxation of ProductMIO. The following proposition formalizes the claim regarding the strength of the two formulations.
@@
Finally, while ProductMIO appears as a tightened reformulation of SplitMIO, its structure connects to another type of formulation in the literature. In particular, a stream of papers in the mixed-integer optimization community (Vielma et al. 2010, Vielma and Nemhauser 2011, Huchette and Vielma 2019) has considered a general approach for deriving small and strong formulations of disjunctive constraints using independent branching schemes; we briefly review the most general such approach from Huchette and Vielma (2019) and showcase its connection to ProductMIO. In this approach, one has a finite ground set $J$, and is interested in optimizing over a particular subset of the (|J|1)(|J|-1)-dimensional unit simplex over $J$, $\Delta^{J}=\{\boldsymbol{\lambda}\in\mathbb{R}^{J}\mid\sum_{j\in J}\lambda_{j}=1;\boldsymbol{\lambda}\geq\mathbf{0}\}$. The specific subset that we are interested in is called a combinatorial disjunctive constraint (CDC), and is given by $\text{CDC}(\mathcal{S})=\bigcup_{S\in\mathcal{S}}Q(S)$, where $\mathcal{S}$ is a finite collection of subsets of $J$ and $Q(S)=\{\boldsymbol{\lambda}\in\Delta\mid\lambda_{j}\leq 0\ \text{for}\ j\in J\setminus S\}$ for any $S\subseteq J$. This approach is very general: for example, by associating each $j$ with a point $\mathbf{x}^{j}$ in $\mathbb{R}^{n}$, one can use $\text{CDC}(\mathcal{S})$ to model an optimization problem over a union of polyhedra, where each polyhedron is the convex hull of a collection of vertices in $S\in\mathcal{S}$.
A $k$-way independent branching scheme of depth $t$ is a representation of $\text{CDC}(\mathcal{S})$ as a sequence of $t$ choices between $k$ alternatives. Specifically, we have $\text{CDC}(\mathcal{S})=\bigcap_{m=1}^{t}\bigcup_{i=1}^{k}Q({L^{i}_{m}})$, where ${L^{i}_{m}}\subseteq J$. In the special case that $k=2$, we can write $\text{CDC}(\mathcal{S})=\cap_{m=1}^{t}(L_{{m}}\cup R_{{m}})$ where $L_{{m}},R_{{m}}\subseteq J$. This representation is known as a pairwise independent branching scheme and the constraints of the corresponding MIO can be written simply as
@@
This particular special case is important because it is always integral (see Theorem 1 of Vielma et al. 2010). Moreover, we can see that ProductMIO bears a strong resemblance to formulation (8). Constraints (8a) and (8b) correspond to constraints (7c) and (7d), respectively. In terms of variables, the $\lambda_{j}$ and $z_{m}$ variables in formulation (8) correspond to the $y_{t,\ell}$ and $x_{i}$ variables in ProductMIO, respectively.
@@
3.5 Comparative Insights on the SplitMIO and ProductMIO Formulations
Another perspective on the difference between the SplitMIO and ProductMIO formulations is to analyze their behavior when $|F|=1$. As shown in Section 9, the feasible region $\mathcal{F}_{\textsc{SplitMIO}}$ of SplitMIO is integral if and only if each product appears at most once in the tree. By contrast, when $|F|=1$, the feasible region $\mathcal{F}_{\textsc{ProductMIO}}$ of ProductMIO is always integral (see Propositions 9.1 and 9.2). This analysis is meaningful for two reasons. First, it shows that a fundamental distinction between the two formulations lies in whether products may appear repeatedly within an individual tree, which is consistent with the way we motivate the ProductMIO formulation in Section 3.4. Second, in the integer programming literature, it is often observed that, among formulations of comparable size for the same problem, those that exhibit desirable properties in simplified settings tend to scale more effectively in the general case (Vielma 2015). For example, Anderson et al. (2020) develop a strong mixed-integer programming formulation for neural networks and justify its strength by analyzing the single-neuron case, even though these properties do not extend to multiple neurons. Our analysis in Section 9 follows a similar logic. Although the integrality results do not extend to forests with multiple trees, they provide valuable insight into the structural differences between the two formulations and deepen our understanding of their relative strength.
For computational and practical purposes, as we illustrate in Section 5.2, ProductMIO is generally preferable to SplitMIO when the assortment problem fits into computer memory and both formulations can be solved directly using commercial solvers such as Gurobi (Gurobi Optimization, LLC 2024). In our numerical experiments, this corresponds to the regime where the number of products is moderate (up to $n=500$). In this setting, ProductMIO clearly outperforms SplitMIO when the trees are highly structured and products are frequently repeated across splits, and shows a smaller but still noticeable edge when the trees are less structured. By contrast, when the problem size increases further (up to $n=3000$), SplitMIO becomes preferable due to the availability of a tailored Benders decomposition method developed in Section 4. In particular, for SplitMIO, Benders cuts can be generated efficiently for both integral and fractional outer decisions $\mathbf{x}$, using closed-form expressions in the integral case and our proposed primal-dual greedy algorithm in the fractional case. For ProductMIO, however, such efficient cut generation is only available when $\mathbf{x}$ is integral. This gives SplitMIO a distinct advantage in large-scale settings during the relaxation phase of Benders decomposition. We demonstrate this advantage in Section 12.5.